OpenStudy (anonymous):

56 divided by [ 48 divided by ( 84 divided by 7) ]

OpenStudy (anonymous):

... anyone

OpenStudy (anonymous):

56 / ( 48 / ( 84 / 7 ) ) = 56 / ( 48 / (12) ) = 56 / (4) = 14

OpenStudy (anonymous):

thanks you think u can help me with some more

OpenStudy (anonymous):

post them on the left :)

OpenStudy (anonymous):

:) thanks!